I have 60% concentrate, I need to mix a 1000 gallons of water to a .60 solution strength. how much concentrate do I need to use?
Answer by Alan3354(69443)

I have 60% concentrate, I need to mix a 1000 gallons of water to a .60 solution strength. how much concentrate do I need to use?
-------------
.60 is 60%
If you mean 0.6%, then
1000*0 + 60*C = (1000 + C)*0.6
60C = 600 + 0.6C
59.4C = 600
C = 600/59.4
C =~ 10.1 gallons of the 60% concentrate.
------------
That will give 1010.1 gallons of 0.6% solution